# Assessing realistic binding energies of some essential interstellar radicals with amorphous solid water:

### v1

@@
Assessing realistic binding energies of some essential interstellar radicals with amorphous solid water:
In the absence of laboratory data, state-of-the-art quantum chemical approaches can provide estimates on the binding energy (BE) of interstellar species with grains. Without BE values, contemporary astrochemical models are compelled to utilize wild guesses, often delivering misleading information. Here, we employ a fully quantum chemical approach to estimate the BE of 7 diatomic radicals – CH, NH, OH, SH, CN, NS, and NO, playing a crucial role in shaping the interstellar chemical composition, using a suitable amorphous solid water model as a substrate since water is the principal constituent of interstellar ice in dense and shielded regions. While the BEs are compatible with physisorption, the binding of CH in some sites shows chemisorption, in which a chemical bond to an oxygen atom of a water molecule is formed. While no structural change can be observed for the CN radical, it is believed that the formation of a hemibonded system between the outer layer of the water cluster and the radical is the reason for the unusually large BE in one of the binding sites considered in our study. A significantly lower BE for NO, consistent with recent calculations, is obtained, which is useful in explaining the recently observed HONO/NH2OH and HONO/HNO ratio in the low-mass hot corino, IRAS 16293-2422B, with chemical models.
@@
Currently, more than 300 molecules have been detected in the interstellar medium or circumstellar shells111https://cdms.astro.uni-koeln.de/classic/molecules (McGuire, 2022). Interstellar grains act as catalysts in shaping the chemical composition of the interstellar clouds (Hasegawa et al., 1992; Das et al., 2008a, 2010, 2016, 2019; Das & Chakrabarti, 2011; Srivastav et al., 2022; Sil et al., 2018, 2021; Ghosh et al., 2022; Mondal et al., 2021). is the most dominant ice component in dense and cold regions of interstellar clouds (Gibb et al., 2004; Boogert et al., 2015). It can account for the % of the ice along most lines of sight (Whittet, 2003). In dense and shielded regions, multiple layers of ice would form either by the direct adsorption of gas phase species or by chemical processing on the grain surface. The proper structure of interstellar ice remains unclear, although it is generally agreed that it is amorphous upon formation (Hama & Watanabe, 2013).
@@
Estimating binding energies (hereafter BEs) of interstellar species is essential to interpret observations and reconstruct our astrochemical history. Several recent studies attempted to provide theoretical BEs of interstellar species. Sil et al. (2017) calculated the BE of most abundant H and on different types of grain surfaces and found that the BE of is always higher than that of H; Wakelam et al. (2017) reported the physisorbed BE of more than interstellar species, including radicals, using a single molecule (monomer) as an approximate amorphous solid water (ASW) ice surface based on quantum chemical calculations. They compared the calculated BEs with experimental values for some molecules. They concluded that the calculated values of the physisorbed BE considering a single molecule are proportional to the experimental values. Das et al. (2018) attempted to improve on this using different sizes of clusters containing up to 6 molecules to calculate the BEs of species, including radicals of astrochemical interest and noted that with increasing cluster size, the calculated BE values start to converge toward the experimentally obtained values. Shimonishi et al. (2018) introduced molecular dynamic (MD) simulations for the first time to properly describe the ASW cluster of 20 molecules and their interactions with atomic C, N, and O using quantum chemistry. Ferrero et al. (2020) employed density functional theory (DFT) using B3LYP-D3(BJ) (for closed-shell molecules) and M06-2X (for radical species) functionals and Ahlrichs’ triple-zeta quality VTZ basis set, supplemented with a double set of polarization functions (A-VTZ*) to calculate the BEs of molecules and radicals (consisting of up to binding sites per each), considering both crystalline and amorphous periodic ice (consisting of ) models. They found good agreement with the previously mentioned studies. In a similar vein, Perrero et al. (2022) and Martínez-Bachs et al. (2024) investigated the BEs of various S- and N-bearing molecules and radicals adsorbed on similar water ice models using similar DFT and higher-level composite methodological approaches. Tinacci et al. (2022) developed a computational framework to simulate the formation of interstellar ice mantles through H2O accretion. They used the ONIOM approach to compute the BE distribution of NH3 considering the ASW ice model composed of 200 H2O molecules. Their results highlighted the complexity of adsorption sites and the potential for multiple BE regimes. In a later study, Tinacci et al. (2023) calculated the distribution of the H2O BE on the same ASW grain model and integrated it into a protoplanetary disk model. They assessed its effect on the position of the water snow line and estimated the water content of potential planetesimals at different distances from the disk center. Bariosco et al. (2024) expanded on this by investigating the distribution of the H2S BE on the same ASW ice surfaces, emphasizing the importance of considering a multitude of binding sites. Duflot et al. (2021) applied the ONIOM high-quantum-mechanic (QM) part and a low-QM part hybrid method to calculate the BEs for a series of atoms, small molecules, and radicals on both crystalline and amorphous ice represented by a cluster of about molecules obtained via classical MD and electronic structure methods. Sameera et al. (2021) used the ONIOM QM/MM method for binding sites of the radical BEs calculation on both crystalline and amorphous ice clusters containing molecules ( in the QM and in the MM region). Enrique-Romero et al. (2022) computed the BEs of radicals using DFT calculations on two ice water cluster models made of and molecules. Bovolenta et al. (2022) computed BE distributions of interstellar molecules and radicals on an amorphized set of water clusters of molecules each, containing unique binding sites. Piacentino & Öberg (2022) published the BEs of P-bearing species on molecule clusters using DFT (M06-2X/aug-cc-pVDZ level of theory). Hendrix et al. (2023) reported computational searches for optimal structures and benchmark binding and condensation energies for sets of neutral, radical, cationic, and anionic molecules of astrochemical relevance with clusters of N molecules. They demonstrated excellent agreement (with discrepancies less than K) between the DFT (B97X-V/def2-svpd and B97X-V/def2-qzvppd level of theories) and coupled-cluster (CCSD(T)/aug-cc-pVTZ level of theory) BEs, which validates the DFT protocols. This could make the DFT level of theory the primary method for studying larger clusters.
@@
The aforementioned studies prevalently focus on the closed-shell species, where all the electrons are paired. In contrast, open-shell species (like radicals) pose technical challenges due to convergence issues in self-consistent field (SCF) procedures. The energy surface of radicals is more complex, with multiple local minima, making it harder to find the global minimum and the unpaired electrons result in a high electron correlation effect, requiring a more sophisticated treatment. Experimentally determining the desorption data of unstable radicals is more challenging due to difficulties in their preparation and deposition, and their tendency to react before desorbing. The chemical energy released upon reaction can also trigger a desorption event (Minissale et al., 2022). Therefore, focusing on stable species is more common due to the difficulties in handling molecular radicals experimentally.
@@
The present work provides the BEs on various amorphous water ices for some open-shell neutral radicals that can be directly implemented into astrochemical models. In the past, there was a prevailing belief that radicals only diffused at relatively high temperatures (¿30 K; Garrod et al., 2008; Das et al., 2015). However, contemporary research suggests that complex organic molecules (COMs) can be generated through non-diffusive chemistry (Shingledecker et al., 2019; Jin & Garrod, 2020; Ioppolo et al., 2021). Since water is the major constituent of a grain mantle in molecular clouds (Keane et al., 2001; Das et al., 2010; Das & Chakrabarti, 2011; Das et al., 2016), we perform BE computations of the radicals CH, NH, OH, SH, CN, NS, and NO considering 9 different ASW clusters constructed with 20 molecules (). A detailed comparison is also made between the calculated BEs and previously available experimental or theoretically obtained BE values.
@@ 2 Computational details and methodology
The BE is usually seen as a local property arising from the electronic interaction between the substrate (grain surface or adsorbent) and the species deposited on it (adsorbate). For a bounded adsorbate, the BE is a positive quantity. It is defined as the difference in electronic optimized energy between the separated radical () and substrate () and the adsorbed radical on the substrate (), i.e.,
@@
For our BE investigation, we consider nine different structures of ASW clusters (see LABEL:tab:opt_water) as the adsorbents. Each ASW ice model was created by cutting out a block from a large unit cell of crystalline ice. They are heated to 300 K using molecular dynamics (MD) annealing calculations based on classical force fields. To prevent water molecules from escaping the cluster, 20 water molecules using the TIP3P model are selected, and droplet simulations are run with a spherical restriction of 20 . Following 100 ps simulations at 300 K, the nine distinct structures are rapidly cooled to a low temperature of 10 K (Shimonishi et al., 2018). These ASW structures are a good compromise between accuracy and computational time/cost for several radicals. Also, ice in astrophysical environments, such as icy satellites, comets, planetary rings, interstellar grains, etc., are mainly amorphous, and the main component of the ice mantle is (Boogert et al., 2015). Moreover, some recent studies (Germain et al., 2022; Tinacci et al., 2022) suggest that a large number (at least a few hundred) of molecules in the ASW cluster could mimic the interstellar ice mantle. However, these types of ASW are too large to handle the complete quantum chemical treatment (geometric optimization and harmonic frequency evaluation). Considering a large number of binding sites, cluster would be a reasonable choice to study the BE. The smaller size of ASW clusters results in the formation of ?under-coordinated? water molecules at their surface, thus pushing up their affinity toward adsorbates and higher BE values. The larger ASW clusters, on the contrary, tend to maximize the number of internal H-bonds. This gives them a more spherical shape and thus has fewer available strong adsorption sites for the adsorbates (Martínez-Bachs et al., 2024) due to a reduction in edges and/or cavities in the model.
@@
All the quantum chemical calculations are performed using the Gaussian 09 and Gaussian 16 suite of programs (Frisch et al., 2009, 2013, 2016, 2019). The considered nine different structures of clusters are fully optimized based on DFT calculations using B97X-D functional (Chai & Head-Gordon, 2008) approach, in which the van der Waals (vdW) interaction (dispersion correction) is empirically incorporated. The choice of B97X-D was motivated by its demonstrated accuracy in describing systems with unpaired electrons, such as radicals. The range-separated hybrid nature of this functional enables a balanced treatment of both electrostatic and dispersion interactions, which are crucial for accurately modeling radical-water complexes. While benchmarking against higher levels of theory is beyond the scope of this work, we acknowledge its significance and intend to address this in future studies. A valence triple- plus polarization and diffuse functions, namely 6-311+G(d,p) basis set, is adopted for the geometry optimization and the BE evaluation. A fully optimized electronic ground state structure, i.e., the local minimum, is verified to be a stationary point (having non-negative frequency) by harmonic vibrational frequency analysis. All the BEs are evaluated with the harmonic vibrational zero-point energy (ZPE) corrections to consider the molecular vibration at low temperatures. Unless otherwise stated, all the BEs reported or compared in this work are the ZPE-corrected BEs. The basis set superposition error (BSSE) is taken care of by the counterpoise correction (CPC) method (Boys & Bernardi, 1970) to get an idea of the BSSE dependency of the system (see 3.2).
@@
We use clusters to simulate the interstellar ASW surface and investigate the influence of H-bond cooperativity on radical BEs as well as the effect of the BE on binding site variation (Bovolenta et al., 2020; Ferrero et al., 2020; Duflot et al., 2021; Bovolenta et al., 2022). The accessible radical binding sites are chosen based on the available dangling-H (-H) positions on ASW cluster surfaces. The BE is controlled by electrostatic attractions, orbital interactions, and Pauli repulsion. As a result, each radical species can have a range of possible BEs (Minissale et al., 2022). As per the information provided in LABEL:tab:opt_water, it is seen that a total of 75 -H sites are available; these sites are distributed among ASW cluster 1 to 9, having 7, 7, 9, 10, 8, 9, 8, 9, and 8 -H sites, respectively. Using a weak H-like bond, geometry optimizations of the complex structures are performed after manually connecting the radical with the -H. During this weak bond connection, we always prefer the largest differences in electronegativity values (see Table 1) between two binding elements of the atom in the radical and -H atom in the cluster. The electronegativity of each atom contributes to the formation of molecular dipoles, which can be strategically utilized to construct the initial adsorption geometries manually. By following the principle of electrostatic complementarity, we always prefer to connect the -H atom of the cluster with the atom in the radical that has the higher electronegativity. This approach leverages the natural tendency of atoms with higher electronegativity to attract electrons or electron density toward themselves, resulting in stronger dipole interactions and more stable adsorption configurations. The electronic ground state spin multiplicity of each of the 7 radicals was checked and noted in Table 2. All the considered radicals are doublets (with one unpaired electron) except NH, which is in the triplet (with two unpaired electrons of the same spin) state. The coordinates of all the representative structures are publicly available in a GitHub repository222https://github.com/milansil/ASW_radicals_data.git.
@@
It is to be noted that the optimized energy and the calculated BEs obtained from the Gaussian program are not always accurate due to stationary point convergence in frequency domain calculations. This is because the root-mean-square (RMS) and maximum force are calculated analytically in the self-consistent field (SCF) and frequency domain. However, SCF numerically solves the maximum and RMS displacement; on the contrary, the post-SCF (frequency) solves both of them analytically. Hence, some deviation of K can be observed. Due to time constraints and high computational cost, the values are taken even when only the maximum and RMS forces are being converged in the frequency domain, and the frequencies are all positive.
3 Computational results and discussions

### v2

@@
Assessing realistic binding energies of some essential interstellar radicals with amorphous solid water
In the absence of laboratory data, state-of-the-art quantum chemical approaches can provide estimates of the binding energy (BE) of interstellar species with grains. Without BE values, contemporary astrochemical models are compelled to utilize wild guesses, often delivering misleading information. Here, we employed a fully quantum chemical approach to estimate the BE of seven diatomic radicals – CH, NH, OH, SH, CN, NS, and NO – that play a crucial role in shaping the interstellar chemical composition, using a suitable amorphous solid water model as a substrate since water is the principal constituent of interstellar ice in dense and shielded regions. While the BEs are compatible with physisorption, the binding of CH in some sites shows chemisorption, in which a chemical bond to an oxygen atom of a water molecule is formed. While no structural change has been observed for the CN radical, it is believed that the formation of a hemibonded system between the outer layer of the water cluster and the radical is the reason for the unusually large BE in one of the binding sites considered in our study. A significantly lower BE for NO, consistent with recent calculations, is obtained, which helps explain the recently observed HONO/NH2OH and HONO/HNO ratios in the low-mass hot corino IRAS 16293–2422 B with chemical models.
@@
Currently, more than 300 molecules have been detected in the interstellar medium or circumstellar shells111https://cdms.astro.uni-koeln.de/classic/molecules (McGuire, 2022). Interstellar grains act as catalysts in shaping the chemical composition of interstellar clouds (Hasegawa et al., 1992; Das et al., 2008a, 2010, 2016, 2019, 2021; Das & Chakrabarti, 2011; Srivastav et al., 2022; Sil et al., 2018, 2021; Ghosh et al., 2022; Mondal et al., 2021). is the most dominant ice component in dense and cold regions of interstellar clouds (Gibb et al., 2004; Boogert et al., 2015). It can account for % of the ice along most lines of sight (Whittet, 2003). In dense and shielded regions, multiple layers of ice would form either via the direct adsorption of gas-phase species or via chemical processing on the grain surface. The proper structure of interstellar ice remains unclear, although it is generally agreed that it is amorphous upon formation (Hama & Watanabe, 2013).
@@
Estimating the binding energies (BEs) of interstellar species is essential for interpreting observations and reconstructing our astrochemical history. Several recent studies attempted to provide theoretical BEs of interstellar species. Sil et al. (2017) calculated the BE of H and (the most abundant species) on different types of grain surfaces and found that the BE of is always higher than that of H; Wakelam et al. (2017, see their Table 2) reported the BEs of interstellar species, including radicals, using a single molecule (monomer) as an approximate amorphous solid water (ASW) ice surface based on quantum chemical calculations. They compared the calculated BEs with experimental values for some molecules. They conclude that the calculated values of the physisorbed BE considering a single molecule as substrate are proportional to the experimental values. Das et al. (2018) attempted to improve on this using different sizes of clusters containing up to six molecules to calculate the BEs of species, including radicals of astrochemical interest, and noted that with increasing cluster size, the calculated BE values start to converge toward the experimentally obtained values. Shimonishi et al. (2018) introduced molecular dynamic (MD) simulations to describe the ASW cluster of 20 molecules and their interactions with atomic C, N, and O using quantum chemistry. Ferrero et al. (2020) employed density functional theory (DFT) using B3LYP-D3(BJ) (for closed-shell molecules) and M06-2X (for radical species) functionals and the Ahlrichs’ triple-zeta quality VTZ basis set, supplemented with a double set of polarization functions (A-VTZ*), to calculate the BEs of molecules and radicals (consisting of up to binding sites each), taking both crystalline and amorphous periodic ice models (consisting of ) into consideration. They find good agreement with the previously mentioned studies. In a similar vein, Perrero et al. (2022) and Martínez-Bachs et al. (2024) investigated the BEs of various S- and N-bearing molecules and radicals adsorbed on similar water ice models using similar DFT and higher-level composite methodological approaches. Tinacci et al. (2022) developed a computational framework to simulate the formation of interstellar ice mantles through H2O accretion. They used the ONIOM (Our own N-layered Integrated molecular Orbital and molecular Mechanics) approach to compute the BE distribution of NH3 for an ASW ice model composed of 200 H2O molecules. Their results highlight the complexity of adsorption sites and the potential for multiple BE regimes. In a later study, Tinacci et al. (2023) calculated the distribution of the H2O BE on the same ASW grain model and integrated it into a protoplanetary disk model. They assessed its effect on the position of the water snow line and estimated the water content of potential planetesimals at different distances from the disk center. Bariosco et al. (2024) expanded on this by investigating the distribution of the H2S BE on the same ASW ice surfaces, emphasizing the importance of considering a multitude of binding sites. Duflot et al. (2021) applied the ONIOM high-quantum-mechanic (QM) part and a low-QM part hybrid method to calculate the BEs for a series of atoms, small molecules, and radicals on both crystalline and amorphous ice represented by a cluster of about molecules obtained via classical MD and electronic structure methods. Sameera et al. (2021) used the ONIOM QM/MM (MM denotes molecular mechanics) method to calculate the BE for ten binding sites of the radical on both crystalline and amorphous ice clusters containing molecules ( in the QM and in the MM region). Enrique-Romero et al. (2022) computed the BEs of radicals using DFT calculations on two ice water cluster models made of and molecules. Bovolenta et al. (2022) computed BE distributions of interstellar molecules and radicals on an amorphized set of water clusters of molecules each, for a total of unique binding sites. Piacentino & Öberg (2022) published the BEs of ten P-bearing species on molecule clusters using DFT (M06-2X/aug-cc-pVDZ level of theory). Hendrix et al. (2023) reported computational searches for optimal structures, benchmark binding, and condensation energies for sets of neutral, radical, cationic, and anionic molecules of astrochemical relevance with clusters of N molecules. They demonstrated excellent agreement (with discrepancies of less than K) between the DFT (B97X-V/def2-svpd and B97X-V/def2-qzvppd levels of theory) and coupled-cluster (CCSD(T)/aug-cc-pVTZ level of theory) BEs, which validates the DFT protocols. This could make the DFT the primary method for studying larger clusters.
@@
The aforementioned studies predominately focused on the closed-shell species, where all the electrons are paired. In contrast, open-shell species (such as radicals) pose technical challenges due to convergence issues in self-consistent field (SCF) procedures. The energy surface of radicals is more complex, with multiple local minima, making it harder to find the global minimum; additionally, the unpaired electrons result in a high electron correlation effect, requiring a more sophisticated treatment. Experimentally determining the desorption data of unstable radicals is more challenging due to difficulties in their preparation and deposition, and their tendency to react before desorbing. The chemical energy released upon a reaction can also trigger a desorption event (Minissale et al., 2022). Therefore, focusing on stable species is more common due to the difficulties in handling molecular radicals experimentally.
@@
The present work provides the BEs on various amorphous water ices for some open-shell neutral radicals that can be directly implemented into astrochemical models. In the past, there was a prevailing belief that radicals only diffused at relatively high temperatures (¿30 K; Garrod et al., 2008; Das et al., 2015). However, contemporary research suggests that complex organic molecules can be generated through non-diffusive chemistry (Shingledecker et al., 2019; Jin & Garrod, 2020; Ioppolo et al., 2021). Since water is the major constituent of a grain mantle in molecular clouds (Keane et al., 2001; Das et al., 2010; Das & Chakrabarti, 2011; Das et al., 2016), we performed BE computations of the radicals CH, NH, OH, SH, CN, NS, and NO, using nine different ASW clusters constructed with 20 molecules (). We also made a detailed comparison of the calculated BEs and previously available experimental or theoretically obtained BE values.
@@ 2 Computational details and methodology
The BE is usually seen as a local property arising from the electronic interaction between the substrate (grain surface or adsorbent) and the species deposited on it (adsorbate). For a bounded adsorbate, the BE is a positive quantity. It is defined as the difference in electronic optimized energy between the separated radical () plus the substrate () and the adsorbed radical on the substrate ():
@@
For our BE investigation, we considered nine different structures of ASW clusters (see Table A.1, available on Zenodo) as the adsorbents. Each ASW ice model was created by cutting out a block from a large unit cell of crystalline ice. They were heated to 300 K using MD annealing calculations based on classical force fields. To prevent water molecules from escaping the cluster, 20 water molecules were selected using the TIP3P model, and droplet simulations were run with a spherical restriction of 20 . Following 100 ps simulations at 300 K, the nine distinct structures were rapidly cooled to a low temperature of 10 K (Shimonishi et al., 2018). These ASW structures are a good compromise between accuracy and computational time and/or cost for several radicals. Also, ice in astrophysical environments, such as icy satellites, comets, planetary rings, and interstellar grains, is mainly amorphous, and the main component of the ice mantle is (Boogert et al., 2015). Moreover, some recent studies (Germain et al., 2022; Tinacci et al., 2022) suggest that a large number of molecules in the ASW cluster (at least a few hundred) could mimic the interstellar ice mantle. However, these types of ASW are too large to handle the complete quantum chemical treatment (geometric optimization and harmonic frequency evaluation). For a large number of binding sites, the cluster would be a reasonable choice for studying the BE. The smaller size of ASW clusters results in the formation of ?under-coordinated? water molecules at their surface, making them more likely to have adsorbates and higher BE values. The larger ASW clusters, on the contrary, tend to have the maximum number of internal H bonds. This gives them a more spherical shape, and thus they have fewer available strong adsorption sites for the adsorbates (Martínez-Bachs et al., 2024) due to a reduction in edges and/or cavities in the model.
@@
All the quantum chemical calculations were performed using the Gaussian 09 and Gaussian 16 suite of programs (Frisch et al., 2009, 2013, 2016, 2019). The nine different structures of clusters we examined were fully optimized based on DFT calculations using the B97X-D functional approach (Chai & Head-Gordon, 2008), in which the van der Waals interaction (dispersion correction) is empirically incorporated. The choice of B97X-D was motivated by its demonstrated accuracy in describing systems with unpaired electrons, such as radicals. The range-separated hybrid nature of this functional enables a balanced treatment of both electrostatic and dispersion interactions, which are crucial for accurately modeling radical-water complexes. While benchmarking against higher levels of theory is beyond the scope of this work, it is important and we intend to do so in future studies. A valence triple- plus polarization and diffuse functions, namely the 6-311+G(d,p) basis set, was adopted for the geometry optimization and the BE evaluation. We verified that a fully optimized electronic ground-state structure (i.e., the local minimum) is a stationary point (with a nonnegative frequency) via harmonic vibrational frequency analysis. All the BEs were evaluated with the harmonic vibrational zero-point energy corrections to study the molecular vibration at low temperatures. Unless otherwise stated, all the BEs discussed in this work are the zero-point-energy-corrected BEs. We corrected for the basis set superposition error (BSSE) using the counterpoise correction (CPC) method (Boys & Bernardi, 1970) to get an idea of the BSSE dependence of the system (see Sect. 3.2).
@@
We used clusters to simulate the interstellar ASW surface and investigate the influence of H-bond cooperativity on radical BEs as well as the effect of the BE on binding site variation (Bovolenta et al., 2020; Ferrero et al., 2020; Duflot et al., 2021; Bovolenta et al., 2022). The accessible radical binding sites were chosen based on the available dangling-H (-H) positions on ASW cluster surfaces. The BE is controlled by electrostatic attractions, orbital interactions, and Pauli repulsion. As a result, each radical species can have a range of possible BEs (Minissale et al., 2022). As per the information provided in Table A.1 (available on Zenodo), a total of 75 -H sites are available; these sites are distributed among ASW clusters 1 to 9 and have 7, 7, 9, 10, 8, 9, 8, 9, and 8 -H sites, respectively. Using a weak H-like bond, geometry optimizations of the complex structures were performed after manually connecting the radical with the -H. During this weak bond connection, we always took the largest differences in electronegativity values (see Table 1) between two binding elements of the atom in the radical and -H atom in the cluster. The electronegativity of each atom contributes to the formation of molecular dipoles, which can be strategically utilized to construct the initial adsorption geometries manually. By following the principle of electrostatic complementarity, we always connected the -H atom of the cluster with the atom in the radical that has the higher electronegativity. This approach leverages the natural tendency of atoms with higher electronegativity to attract electrons or electron density toward themselves, resulting in stronger dipole interactions and more stable adsorption configurations. The electronic ground-state spin multiplicity of each of the seven radicals was checked and is noted in Table 2. All the considered radicals are doublets (with one unpaired electron) except NH, which is in the triplet state (with two unpaired electrons of the same spin). The coordinates of all the representative structures are publicly available in a GitHub repository222https://github.com/milansil/ASW_radicals_data.git.
@@
It is to be noted that the optimized energy and the calculated BEs obtained from the Gaussian program are not always accurate, due to stationary point convergence in frequency domain calculations. This is because the root mean square (RMS) and the maximum force are calculated analytically in the SCF and frequency domain. However, SCF numerically solves the maximum and RMS displacement; on the contrary, the post-SCF (frequency) solves both of them analytically. Hence, some deviation of K can be observed. Due to time constraints and high computational costs, the values were taken even when only the maximum and RMS forces were converged in the frequency domain, and the frequencies are all positive.
3 Computational results and discussion